OpenStudy (anonymous):

Surface Integral z=5-sqrt(x^2+y^2) above the z=0 plane, integrate (x^2+y^2+5z^2)

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

\[\int\limits_{0}^{5}\int\limits_{0}^{\sqrt{25-y^2}} (x^2+y^2+5(5-\sqrt{x^2+y^2})^2)*\sqrt((-x/(x^2+y^2)^2 + (-y/(x^2+y^2))^2 +1))\]

OpenStudy (anonymous):

wow that is messy to look at. but that's what i got

OpenStudy (ipwnbunnies):

Lol, I'll give a hint. Do the parametricizing of the surface. Do all the cross product stuff. Find the region of the cone, which is easy. Try using spherical coordinates for your integration.

OpenStudy (ipwnbunnies):

Ugh, I mean polar coordinates. This is a double integral.

OpenStudy (anonymous):

instead of doing the cross product, I just use the formula sqrt(dz/dx^2 + dz/dy^2 +1)

OpenStudy (anonymous):

So I turned z=x^2+y^2+5z^2 into z=x^2+y^2+5(5-sqrt(x^2+y^2))^2

OpenStudy (anonymous):

from that, I was able to get the bounds of 0 to 5 for y and 0 to sqrt(25-y^2) for x

OpenStudy (kainui):

I am definitely on board with @iPwnBunnies, find out what the cone part looks like and it'll be a lot simpler to look at and write down.

OpenStudy (anonymous):

I am awful with polar coordinates

OpenStudy (ipwnbunnies):

Lol, I find it easy to convert. We already know the cone has a circular region with radius 5. That should be plenty...I am now out. Peace Kainui and Mgb.

OpenStudy (anonymous):

Thanks @iPwnBunnies much appreciate the help

OpenStudy (anonymous):

so I got, z=r^2+5(5-r)

OpenStudy (anonymous):

i mean for the function after plugging z in

OpenStudy (anonymous):

\[\int\limits_{0}^{2*\pi}\int\limits_{0}^{5} (r^2-5r+25) *r \] dr dtheta

OpenStudy (anonymous):

but im missing the sqrt(dz/dx^2 + dz/dy^2 +1) part still

OpenStudy (anonymous):

i got sqrt(2) for that part

OpenStudy (kainui):

Well instead of using that, I would use the cross product version. That I believe is specifically rectangular coordinates and won't work here.

OpenStudy (anonymous):

is the other part correct?

OpenStudy (kainui):

I'm sort of not following your way of doing it to be honest, I'm not sure it's correct until you get an answer in the end for me to compare it to.

OpenStudy (kainui):

I'm using:\[T(r, \theta) = <x,y, 5-\sqrt{x^2+y^2}>\]\[T(r, \theta )=<r \cos \theta, r \sin \theta, 5-r>\]

OpenStudy (anonymous):

ok so then, \[for r, we get <\cos \theta, \sin \theta,-1>\]

OpenStudy (anonymous):

for theta, \[<-rsin \theta,rcos \theta,0>\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

then cross those?

OpenStudy (kainui):

You got it. =)

OpenStudy (anonymous):

alright so we get \[(r \cos(\theta), rsin(\theta),rcos^2(\theta)+rsin^2(\theta))\]

OpenStudy (anonymous):

or <rcos(t), rsin(t), r> right?

OpenStudy (kainui):

Yes, you got it. I'm afraid I have to leave, I've been distracted this whole time but I've got to go I'm sorry!

OpenStudy (anonymous):

the magnetitude would be\[\sqrt((r \cos(x))^2+(r \sin(x))^2+(r)^2)\]

OpenStudy (anonymous):

aww ok thanks for the help

OpenStudy (anonymous):

@mathmale help me! haha